OpenStudy (sloppycanada):

The difference of two numbers is 4. 5 times the smaller number is four more than three times the larger. Find the two numbers.

OpenStudy (sloppycanada):

I think the answer is -10 and -14

OpenStudy (anonymous):

\[x-y=4.5\\ y=3x+4\]is what i would solve

OpenStudy (anonymous):

which does not give the two solutions you wrote

OpenStudy (anonymous):

oh wait, maybe i read it wrong sorry

OpenStudy (anonymous):

actually i just read it again carefully The difference of two numbers is 4. 5 times the smaller number is four more than three times the larger. does not make any sense to me

OpenStudy (sloppycanada):

It doesn't make sense to me either.

OpenStudy (anonymous):

this is simple here, The difference of two numbers is 4 y-x=4 y=4+x

OpenStudy (anonymous):

5 times the smaller number is four more than three times the larger which equals to 5x=3y+4

OpenStudy (anonymous):

simplify 5x=3(x+4)+4

OpenStudy (mertsj):

let x = bigger number let y = smaller number x-y=4 5y= 3x+4

OpenStudy (anonymous):

@sloppycanada are you getting it so far?

OpenStudy (sloppycanada):

I think so.. would my answers be 8 and 12?

OpenStudy (anonymous):

yes thats correct